OpenStudy (anonymous):

What is the sum of 2.5 × 104 and 3.4 × 104? 0.59 × 10^4 0.59 × 10^8 5.9 × 10^4 5.9 × 10^8

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

\[2.5 \times 10^4+3.4\times 10^4=\left( 2.5+3.4 \right)\times 10^4=?\]

OpenStudy (anonymous):

c